Find smallest number when it is divided by 7,9 and 11 with remainder is 1, 2 and 3 respectively.
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)


92
7*13=91 +1=92
9*10=90 +2=92
11*8=88 +3=92













Answer by Alan3354(69443)

Find smallest number when it is divided by 7,9 and 11 with remainder is 1, 2 and 3 respectively.
----------------------
The smallest positive # is 344